OpenStudy (anonymous):

Solve this system of equations using the addition method. x + y = 8 2x = y - 5

OpenStudy (ivettef365):

x + y = 8 2x - y = -5 --------------- 3x = 3 -- -- 3 3 x = 1 Therefore since x=1, then solve for y 1 + y = 8 y = 8 - 1 y = 7 x = 1 and y = 7

OpenStudy (anonymous):

X=1 and y=7